OpenStudy (anonymous):

What is the equation of a line that is parallel to 4x+6y=10 and passes through (-4, 1)

OpenStudy (anonymous):

Anybody

OpenStudy (anonymous):

\[ 4x + 6y = 10 \quad \to \quad 6y = -4x + 10 \\ y = \frac{-4}{6} + \frac{10}{6} = -\frac{2}{3}x + \frac{5}{3} \] by \(y = mx + h\) we get \(m = -\frac{2}{3}\). This means that the slope of this straight line function is \(-\frac{2}{3}\). Now, we need a function in form of \(y = -\frac{2}{3}x + C \) that would go through (-4,1) \\ So, we know x and y for our point, let's put it in to find C: \[ x = -4 \quad \quad \quad y = 1 \\ y = -\frac{2}{3}x + C \quad \to \quad 1 = -\frac{2}{3}(-4) + C \\ 1 = \frac{8}{3} + C \quad \to \quad \frac{3}{3} - \frac{8}{3} = C \\ C = -\frac{5}{3} \] So our function is \[ y = -\frac{2}{3}x - \frac{5}{3} \] Small check: \[ x = -4 \quad \implies \quad y = -\frac{2}{3}(-4) - \frac{5}{3} \\ y = \frac{8}{3} - \frac{5}{3} = \frac{3}{3} = 1 \]